# Sample Complexity of Linear Quadratic Regulator Without Initial Stability

### v2

@@ 5. Sample complexity
$N=\frac{1}{2}\cdot\frac{\log\big{(}\frac{2\|Q_{N}-P^{*}\|_{*}\cdot\kappa_{P^{*}}\cdot\|A_{K}^{*}\|\cdot\|B\|}{\epsilon\cdot\lambda_{\min}(R)}\big{)}}{\log\big{(}\frac{1}{\|A_{K}^{*}\|_{*}}\big{)}}+1,$
@@
$\alpha_{h,t}=\frac{2}{\mu}\frac{1}{t+\theta_{h}}\quad\text{for}\quad\theta_{h}=\max\{2,\frac{2L\xi_{h,4}}{\mu^{2}\widetilde{J}_{h}(K_{h,0})}\},$
@@
outputs a control policy $\widetilde{K}_{0}$ that satisfies $\|\widetilde{K}_{0}-K^{*}\|\leq\varepsilon$ with a probability of at least $1-N\zeta$. Furthermore, if $\varepsilon$ is sufficiently small such that
$\varepsilon<\frac{1-\|A-BK^{*}\|_{*}}{\|B\|},$
then $\widetilde{K}_{0}$ is stabilizing.
The results in Corollary 5.1 provide a rigorous theoretical foundation for Algorithm 1, ensuring it computes a control policy $\widetilde{K}_{0}$ satisfying $\|\widetilde{K}_{0}-K^{*}\|\leq\epsilon$ with high probability. The following corollary formalizes the sample complexity bound of our approach.
@@
$\sum_{h=0}^{N-1}T_{h}=\widetilde{\mathcal{O}}\left(\epsilon^{-2}\right).$
@@
$\sum_{h=0}^{N-1}T^{\prime}_{h}=\widetilde{\mathcal{O}}\left(\max\left\{\epsilon^{-2},\epsilon^{-\left(1+\frac{\log(C_{2})}{2\log{(1/\|A^{*}_{K}\|_{*}})}\right)}\right\}\right),$
@@
This comparison highlights the advantage of our method, which achieves a uniform sample complexity bound of $\widetilde{\mathcal{O}}(\epsilon^{-2})$, independent of problem-specific constants. In contrast, the bound in [20] deteriorates as $C_{2}$ increases, since their second term scales as
$\widetilde{\mathcal{O}}\left(\epsilon^{-\left(1+\frac{\log(C_{2})}{2\log(1/\|A^{*}_{K}\|_{*})}\right)}\right).$
This can be arbitrarily worse than $\widetilde{\mathcal{O}}(\epsilon^{-2})$, leading to much higher sample complexity in some cases.
@@ 6. Simulation Studies
We set $N=\lceil\frac{1}{2}\log\left(\frac{1}{\epsilon}\right)\rceil$, consistent with (31), and in each inner loop, apply the policy gradient (PG) update outlined in Algorithm 1 using a time-varying step-size as suggested in (73). The algorithm is run for twelve different values of $\epsilon$: $\epsilon\in\{10^{-6},10^{-5.5},10^{-5},\dotsc,10^{-0.5}\}$, with the results shown in Figure 1. To account for the inherent randomness in the algorithm, we perform one hundred independent runs for each value of $\varepsilon$ and compute the average sample complexity and policy optimality gap $\|\widetilde{K}_{0}-K^{*}\|$. As seen in Figure 1, the sample complexity exhibits a slope consistent with $\mathcal{O}(\varepsilon^{-0.5})$, visibly outperforming the method in [20], which demonstrates a much steeper slope of approximately $\mathcal{O}(\varepsilon^{-1.5})$.
@@ 7. Conclusion
Beyond these improvements, we introduced a refined outer-loop analysis that significantly enhances error accumulation, leveraging the contraction of the Riemannian distance over the Riccati operator. This ensures that the accumulated error remains linear in the horizon length, leading to a sample complexity bound of $\widetilde{\mathcal{O}}(\epsilon^{-2})$, independent of problem-specific constants, making the method more broadly applicable.
@@
$\overline{P}_{t}:=P^{*}_{t}-P^{*},\quad\overline{R}:=R+B^{\top}P^{*}B,$
@@
$\overline{P}_{t}$ $=\overline{A}^{\top}\overline{P}_{t+1}\overline{A}-\overline{A}^{\top}\overline{P}_{t+1}B(\overline{R}+B^{\top}\overline{P}_{t+1}B)^{-1}B^{\top}\overline{P}_{t+1}\overline{A}$
@@
$\leq\big{[}1+\lambda_{\min}(\overline{P}^{1/2}_{t+1}B\overline{R}^{-1}B^{\top}\overline{P}_{t+1}^{1/2})\big{]}^{-1}\overline{A}^{\top}\overline{P}_{t+1}\overline{A}$
(98) $=:\mu_{t}\overline{A}^{\top}\overline{P}_{t+1}\overline{A},$
@@
Therefore, the convergence is exponential such that $\|\overline{P}_{t}\|_{*}\leq\|\overline{A}\|_{*}^{2(N-t)}\cdot\|\overline{P}_{N}\|_{*}$. As a result, the convergence of $\overline{P}_{t}$ to 00 in spectral norm can be characterized as
@@
where we have used $\kappa_{X}$ to denote the condition number of $X$. That is, to ensure $\|\overline{P}_{1}\|\leq\epsilon$, it suffices to require
(99) $N\geq\frac{1}{2}\cdot\frac{\log\big{(}\frac{\|\overline{P}_{N}\|_{*}\cdot\kappa_{P^{*}}}{\epsilon}\big{)}}{\log\big{(}\frac{1}{\|\overline{A}\|_{*}}\big{)}}+1.$
@@
Substituting $\epsilon$ in (99) with $\frac{\epsilon\cdot\lambda_{\min}(R)}{\|\overline{A}\|\cdot\|B\|}$ completes the proof.
@@
(112) $\varepsilon\leq\frac{1}{N}\min\left\{\frac{a}{2e\|P_{\max}\|},1\right\}.$
@@
(118) $\delta(P_{t^{\prime}-1}^{*},\widetilde{P}_{t^{\prime}-1}^{*})$ =$\overset{\mathrm{(i)}}{=}\delta(\mathcal{R}(P^{*}_{t^{\prime}}),\mathcal{R}(\widetilde{P}_{t^{\prime}}))$
@@
(120) $\delta(P_{t}^{*},\widetilde{P}_{t}^{*})$ $\leq\delta(P_{t+1}^{*},\widetilde{P}_{t+1})$
(121) $\overset{\mathrm{(i)}}{\leq}\delta(P_{t+1}^{*},\widetilde{P}^{*}_{t+1})+\delta(\widetilde{P}_{t+1}^{*},\widetilde{P}_{t+1})$
(122) $\leq\delta(P_{t+2}^{*},\widetilde{P}_{t+2})+\delta(\widetilde{P}_{t+1}^{*},\widetilde{P}_{t+1})$
(123) $\leq\delta(P_{t+2}^{*},\widetilde{P}^{*}_{t+2})+\delta(\widetilde{P}_{t+2}^{*},\widetilde{P}_{t+2})+\delta(\widetilde{P}_{t+1}^{*},\widetilde{P}_{t+1})$
[…]
(125) $\leq\delta(P_{N}^{*},\widetilde{P}_{N})+\sum_{k=1}^{N-t-1}\delta(\widetilde{P}_{t+k}^{*},\widetilde{P}_{t+k})$
(126) =$\overset{\mathrm{(ii)}}{=}\sum_{k=1}^{N-t-1}\delta(\widetilde{P}_{t+k}^{*},\widetilde{P}_{t+k})$
@@
where (i) is due to the triangle inequality of the Riemannian distance [3], (ii) follows from $P_{N}^{*}=\widetilde{P}_{N}=Q_{N}$, and (iii) from (111). We now start the proof of (114) by writing
@@
(129) $\|P^{*}_{t}-\widetilde{P}^{*}_{t}\|$ $\overset{\mathrm{(i)}}{\leq}\|P_{\max}\|e^{\delta(P^{*}_{t},\widetilde{P}^{*}_{t})}\delta(P^{*}_{t},\widetilde{P}^{*}_{t})$
(130) $\overset{\mathrm{(ii)}}{\leq}\|P_{\max}\|e^{\varepsilon N}\varepsilon N$
@@
$=-(\widetilde{K}^{*}_{t})^{\top}B^{\top}\widetilde{P}_{t+1}A-A^{\top}\widetilde{P}_{t+1}B\widetilde{K}_{t}^{*}+(\widetilde{K}^{*}_{t})^{\top}(R+B^{\top}\widetilde{P}_{t+1}B)\widetilde{K}^{*}_{t}$
@@
=$\overset{\mathrm{(i)}}{=}\big{[}(R+B^{\top}\widetilde{P}_{t+1}B)^{-1}B^{\top}\widetilde{P}_{t+1}A-\widetilde{K}^{*}_{t}\big{]}^{\top}(R+B^{\top}\widetilde{P}_{t+1}B)\big{[}(R+B^{\top}\widetilde{P}_{t+1}B)^{-1}B^{\top}\widetilde{P}_{t+1}A-\widetilde{K}^{*}_{t}\big{]}$
@@
(133) $\|\widetilde{P}^{*}_{t}-\widetilde{P}_{t}\|$ $\leq\|R+B^{\top}(P_{\max}+a)B\|\|\widetilde{K}^{*}_{t}-\widetilde{K}_{t}\|^{2}$
@@
(137) $\|\widetilde{P}^{*}_{t}-\widetilde{P}_{t}\|_{F}$ $\leq\|R+B^{\top}(P_{\max}+aI)B\|\|\widetilde{K}^{*}_{t}-\widetilde{K}_{t}\|_{F}^{2}$
@@
Proof of Theorem 3.2: First, according to Theorem 3.1, our choice of $N$ in (31) ensures that $K^{*}_{0}$ is stabilizing and $\|K^{*}_{0}-K^{*}\|\leq\epsilon/2$. Then, it remains to show that the output $\widetilde{K}_{0}$ satisfies $\|\widetilde{K}_{0}-K^{*}_{0}\|\leq\epsilon/2$.
@@
Next, since we have $\|\widetilde{K}_{0}-\widetilde{K}_{0}^{*}\|\leq\varsigma_{0,\varepsilon}=\epsilon/4$, it suffices to show $\|\widetilde{K}^{*}_{0}-K^{*}_{0}\|\leq\epsilon/4$ to fulfill $\|\widetilde{K}_{0}-K^{*}_{0}\|\leq\epsilon/2$. Then, by (143), in order to satisfy $\|\widetilde{K}^{*}_{0}-K^{*}_{0}\|\leq\epsilon/4$, it remains to show
(144) $\|P^{*}_{1}-\widetilde{P}_{1}\|\leq\frac{\epsilon}{4C_{1}}.$
@@
(145) $\varepsilon=\frac{1}{N}\min\left\{\frac{\varepsilon}{8eC_{1}\|P_{\max}\|},\frac{a}{2e\|P_{\max}\|},1\right\},$
@@
(157) $\|P^{*}_{1}-\widetilde{P}^{*}_{1}\|$ $\overset{\mathrm{(i)}}{\leq}\|P_{\max}\|e^{\delta(P^{*}_{1},\widetilde{P}^{*}_{1})}\delta(P^{*}_{1},\widetilde{P}^{*}_{1})$
(158) $\overset{\mathrm{(ii)}}{\leq}\|P_{\max}\|e^{\varepsilon N}\varepsilon N$
@@
(161) $\overset{\mathrm{(i)}}{\leq}\|R+B^{\top}(P_{\max}+aI)B\|(\varsigma_{1,\varepsilon})^{2}$
@@
thereby establishing (144) and concluding the proof of Theorem 3.2. $\diamond$

### v3

@@ 5. Sample complexity
$N=\frac{1}{2}\cdot\frac{\log\big(\frac{2\|Q_{N}-P^{*}\|_{*}\cdot\kappa_{P^{*}}\cdot\|A_{K}^{*}\|\cdot\|B\|}{\varepsilon\cdot\lambda_{\min}(R)}\big)}{\log\big(\frac{1}{\|A_{K}^{*}\|_{*}}\big)}+1,$
@@
$\alpha_{h,t}=\frac{2}{\mu}\frac{1}{t+\theta_{h}}\quad\text{for}\quad\theta_{h}=\max\{2,\frac{2L\xi_{h,3}}{\mu^{2}\widetilde{J}_{h}(K_{h,0})}\},$
@@
outputs a control policy $\widetilde{K}_{0}$ that satisfies $\|\widetilde{K}_{0}-K^{*}\|\leq\varepsilon$ with a probability of at least $1-N\zeta$. Additionally, for a sufficiently small $\varepsilon$ such that $\varepsilon<\frac{1-\|A-BK^{*}\|_{*}}{\kappa_{P^{*}}^{1/2}\ \|B\|}$, $\widetilde{K}_{0}$ is stabilizing.
The results in Corollary 5.1 provide a rigorous theoretical foundation for Algorithm 1, ensuring it computes a control policy $\widetilde{K}_{0}$ satisfying $\|\widetilde{K}_{0}-K^{*}\|\leq\varepsilon$ with high probability. The following corollary formalizes the sample complexity bound of our approach.
@@
$\sum_{h=0}^{N-1}T_{h}=\widetilde{\mathcal{O}}\left(\varepsilon^{-2}\right).$
@@
$\sum_{h=0}^{N-1}T^{\prime}_{h}=\widetilde{\mathcal{O}}\left(\max\left\{\varepsilon^{-2},\varepsilon^{-\left(1+\frac{\log(C_{2})}{2\log{(1/\|A^{*}_{K}\|_{*}})}\right)}\right\}\right),$
@@
This comparison highlights the advantage of our method, which achieves a uniform sample complexity bound of $\widetilde{\mathcal{O}}(\varepsilon^{-2})$, independent of problem-specific constants. In contrast, the bound in [22] deteriorates as $C_{2}$ increases, since their second term scales as
$\widetilde{\mathcal{O}}\left(\varepsilon^{-\left(1+\frac{\log(C_{2})}{2\log(1/\|A^{*}_{K}\|_{*})}\right)}\right).$
This can be arbitrarily worse than $\widetilde{\mathcal{O}}(\varepsilon^{-2})$, leading to much higher sample complexity in some cases.
@@ 6. Simulation Studies
We set $N=\lceil\frac{1}{2}\log\left(\frac{1}{\varepsilon}\right)\rceil$, consistent with (31), and in each inner loop, apply the policy gradient (PG) update outlined in Algorithm 1 using a time-varying step-size as suggested in (69). The algorithm is run for twelve different values of $\varepsilon$: $\varepsilon\in\{10^{-6},10^{-5.5},10^{-5},\dotsc,10^{-0.5}\}$, with the results shown in Figure 3. To account for the inherent randomness in the algorithm, we perform one hundred independent runs for each value of $\varepsilon$ and compute the average sample complexity and policy optimality gap $\|\widetilde{K}_{0}-K^{*}\|$. As seen in Figure 3, the sample complexity exhibits a slope consistent with $\mathcal{O}(\varepsilon^{-0.5})$, visibly outperforming the method in [22], which, as shown in [22, Figure 2], demonstrates a much steeper slope of approximately $\mathcal{O}(\varepsilon^{-1.5})$.
@@ 7. Conclusion
Beyond these improvements, we introduced a refined outer-loop analysis that significantly enhances error accumulation, leveraging the contraction of the Riccati operator under the Riemannian distance. This ensures that the accumulated error remains linear in the horizon length, leading to a sample complexity bound of $\widetilde{\mathcal{O}}(\varepsilon^{-2})$, independent of problem-specific constants, making the method more broadly applicable.
@@
$\overline{P}_{t}:=P^{*}_{t}-P^{*}\hskip-1.99997pt,\quad\overline{R}:=R+B^{\top}P^{*}B,$
@@
$\overline{P}_{t}$ $=\overline{A}^{\top}\overline{P}_{t+1}\overline{A}-\overline{A}^{\top}\hskip-1.00006pt\overline{P}_{t+1}B(\overline{R}+B^{\hskip-0.81949pt\top}\hskip-1.00006pt\overline{P}_{t+1}B)^{\hskip-0.81949pt-1}\hskip-1.00006ptB^{\hskip-0.81949pt\top}\hskip-1.00006pt\overline{P}_{t+1}\overline{A}$
@@
$\leq\hskip-1.00006pt\big[1\hskip-1.49994pt+\hskip-1.49994pt\lambda_{\min}(\overline{P}^{1/2}_{t+1}B\overline{R}^{-1}B^{\top}\overline{P}_{t+1}^{1/2})\big]^{-1}\hskip-1.00006pt\overline{A}^{\top}\overline{P}_{t+1}\overline{A}$
(94) $=:\hskip-1.00006pt\mu_{t}\overline{A}^{\top}\hskip-1.99997pt\overline{P}_{t+1}\overline{A},$
@@
Therefore, the convergence is exponential such that $\|\overline{P}_{t}\|_{*}\leq\|\overline{A}\|_{*}^{2(N-t)}\cdot\|\overline{P}_{N}\|_{*}$. As a result, the convergence of $\overline{P}_{t}$ to 0 in spectral norm can be characterized as
@@
where we have used $\kappa_{X}$ to denote the condition number of $X$. That is, to ensure $\|\overline{P}_{1}\|\leq\varepsilon$, it suffices to require
(95) $N\geq\frac{1}{2}\cdot\frac{\log\big(\frac{\|\overline{P}_{N}\|_{*}\cdot\kappa_{P^{*}}}{\varepsilon}\big)}{\log\big(\frac{1}{\|\overline{A}\|_{*}}\big)}+1.$
@@
Substituting $\varepsilon$ in (95) with $\frac{\varepsilon\cdot\lambda_{\min}(R)}{\|\overline{A}\|\cdot\|B\|}$ completes the proof.
@@
(108) $\varepsilon\leq\frac{1}{N}\min\left\{\frac{a}{2e\|Q_{N}\|},1\right\}.$
@@
(114) $\delta(P_{t^{\prime}-1}^{*},\widetilde{P}_{t^{\prime}-1}^{*})$ $\overset{\mathrm{(i)}}{=}\delta(\mathcal{R}(P^{*}_{t^{\prime}}),\mathcal{R}(\widetilde{P}_{t^{\prime}}))$
@@
(116) $\delta(P^{*}_{t},\widetilde{P}^{*}_{t})$ $\leq\delta(P^{*}_{t+1},\widetilde{P}_{t+1})$
(117) $\overset{\mathrm{(i)}}{\leq}\delta(P^{*}_{t+1},\widetilde{P}^{*}_{t+1})+\delta(\widetilde{P}^{*}_{t+1},\widetilde{P}_{t+1})$
(118) $\leq\delta(P^{*}_{t+2},\widetilde{P}_{t+2})+\delta(\widetilde{P}^{*}_{t+1},\widetilde{P}_{t+1})$
(119) $\leq\delta(P^{*}_{t+2},\widetilde{P}^{*}_{t+2})+\delta(\widetilde{P}^{*}_{t+2},\widetilde{P}_{t+2})+\delta(\widetilde{P}^{*}_{t+1},\widetilde{P}_{t+1})$
(120) $\leq\cdots$
(121) $\leq\delta(P^{*}_{N},\widetilde{P}_{N})+\sum_{k=1}^{N-t-1}\delta(\widetilde{P}^{*}_{t+k},\widetilde{P}_{t+k})$
(122) $\overset{\mathrm{(ii)}}{=}\sum_{k=1}^{N-t-1}\delta(\widetilde{P}^{*}_{t+k},\widetilde{P}_{t+k})$
@@
where (i) is due to the triangle inequality of the Riemannian distance [3], (ii) follows from $P^{*}_{N}=\widetilde{P}_{N}=Q_{N}$, and (iii) from (107). We now start the proof of (110) by writing
@@
(125) $\|P^{*}_{t}-\widetilde{P}^{*}_{t}\|$ $\overset{\mathrm{(i)}}{\leq}\|Q_{N}\|e^{\delta(P^{*}_{t},\widetilde{P}^{*}_{t})}\delta(P^{*}_{t},\widetilde{P}^{*}_{t})$
(126) $\overset{\mathrm{(ii)}}{\leq}\|Q_{N}\|e^{\varepsilon N}\varepsilon N$
@@
$=-(\widetilde{K}^{*}_{t})^{\hskip-0.81949pt\top}\hskip-1.00006ptB^{\hskip-0.81949pt\top}\hskip-1.00006pt\widetilde{P}_{t+1}A\hskip-1.00006pt-\hskip-1.00006ptA^{\hskip-0.81949pt\top}\hskip-1.00006pt\widetilde{P}_{t+1}B\widetilde{K}_{t}^{*}\hskip-1.00006pt+\hskip-1.00006pt(\widetilde{K}^{*}_{t})^{\hskip-0.81949pt\top}\hskip-1.00006pt(R+B^{\hskip-0.81949pt\top}\widetilde{P}_{t+1}B)\widetilde{K}^{*}_{t}$
@@
$\overset{\mathrm{(i)}}{=}\big[(R+B^{\top}\widetilde{P}_{t+1}B)^{-1}B^{\top}\widetilde{P}_{t+1}A-\widetilde{K}^{*}_{t}\big]^{\top}(R+B^{\top}\widetilde{P}_{t+1}B)\big[(R+B^{\top}\widetilde{P}_{t+1}B)^{-1}B^{\top}\widetilde{P}_{t+1}A-\widetilde{K}^{*}_{t}\big]$
@@
(129) $\|\widetilde{P}^{*}_{t}-\widetilde{P}_{t}\|$ $\leq\|R+B^{\top}(Q_{N}+a)B\|\|\widetilde{K}^{*}_{t}-\widetilde{K}_{t}\|^{2}$
@@
(133) $\|\widetilde{P}^{*}_{t}-\widetilde{P}_{t}\|_{F}$ $\leq\|R+B^{\top}(Q_{N}+aI)B\|\|\widetilde{K}^{*}_{t}-\widetilde{K}_{t}\|_{F}^{2}$
@@
Proof of Theorem 3.2: First, according to Theorem 3.1, our choice of $N$ in (31) ensures that $K^{*}_{0}$ is stabilizing and $\|K^{*}_{0}-K^{*}\|\leq\varepsilon/2$. Then, it remains to show that the output $\widetilde{K}_{0}$ satisfies $\|\widetilde{K}_{0}-K^{*}_{0}\|\leq\varepsilon/2$.
@@
Next, since we have $\|\widetilde{K}_{0}-\widetilde{K}_{0}^{*}\|\leq\varsigma_{0,\varepsilon}=\varepsilon/4$, it suffices to show $\|\widetilde{K}^{*}_{0}-K^{*}_{0}\|\leq\varepsilon/4$ to fulfill $\|\widetilde{K}_{0}-K^{*}_{0}\|\leq\varepsilon/2$. Then, by (139), in order to satisfy $\|\widetilde{K}^{*}_{0}-K^{*}_{0}\|\leq\varepsilon/4$, it remains to show
(140) $\|P^{*}_{1}-\widetilde{P}_{1}\|\leq\frac{\varepsilon}{4C_{1}}.$
@@
(141) $\varepsilon=\frac{1}{N}\min\left\{\frac{\varepsilon}{8eC_{1}\|Q_{N}\|},\frac{a}{2e\|Q_{N}\|},1\right\},$
@@
(153) $\|P^{*}_{1}-\widetilde{P}^{*}_{1}\|$ $\overset{\mathrm{(i)}}{\leq}\|Q_{N}\|e^{\delta(P^{*}_{1},\widetilde{P}^{*}_{1})}\delta(P^{*}_{1},\widetilde{P}^{*}_{1})$
(154) $\overset{\mathrm{(ii)}}{\leq}\|Q_{N}\|e^{\varepsilon N}\varepsilon N$
@@
(157) $\overset{\mathrm{(i)}}{\leq}\|R+B^{\top}(Q_{N}+aI)B\|(\varsigma_{1,\varepsilon})^{2}$
@@
thereby establishing (140) and concluding the proof of the first claim of Theorem 3.2.
Lastly, we need to show if $\varepsilon<\frac{1-\|A-BK^{*}\|_{*}}{\kappa_{P^{*}}^{1/2}\ \|B\|}$, then $\widetilde{K}_{0}$ is stabilizing. First, recall that $A-B\widetilde{K}_{0}$ is stable if $\|A-B\widetilde{K}_{0}\|_{*}<1$. Now note that we can write
$\|A-B\widetilde{K}_{0}\|_{*}$ $=\|A-BK^{*}\|_{*}+\|B(K^{*}-\widetilde{K}_{0})\|_{*}$
$=\|A-BK^{*}\|_{*}+\left\|P^{*^{1/2}}B(K^{*}-\widetilde{K}_{0})P^{*^{-1/2}}\right\|$
$\leq\|A-BK^{*}\|_{*}+\|P^{*^{1/2}}\|\|P^{*^{-1/2}}\|\|B\|\|K^{*}-\widetilde{K}_{0}\|$
$\leq\|A-BK^{*}\|_{*}+\kappa_{P^{*}}^{1/2}\ \|B\|\varepsilon$
$<1,$
which verifies the second claim, concluding the proof of Theorem 3.2. $\diamond$